In an atom, an electron is moving with a speed of 600 m/s with an accuracy of 0.005%. Certainity with which the position of the electron can be located is (h = 6.6 × 10–34 kg m2 s–1, mass of electron, em = 9.1 × 10–31 kg) : ?

Grade:12

1 Answers

Sindhu Sundar
4 years ago
velocity, v = 600m/s
accuracy = 0.005%
Thus, uncertainity in velocity = 600*0.005 = 3m/s
According to the uncertainity principle,
Delta(X).Delta(P) >= h/4pi
Delta(P) = m.Delta(V)
Delta(X) = h/4pi.m.Delta(V)
Delta(X) = (6.62*10^-34*7)/(4*22*9.1*10^-31*3)
Delta(X) = 1.9*10^-5